# Models of Set Theory: Extensions and Dead-ends

### v5

@@ 3. TALLER CONSERVATIVE ELEMENTARY EXTENSIONS
If case 1 holds then we shall simply let $X_{n+1,\alpha}=f_{n+1,\alpha}^{-1}\{\beta:\beta<\gamma_{0}\},$ where $\gamma_{0}$ is the first ordinal witnessing the veracity of case 1.
If case 2 holds then we can define a class $S=\{s_{\alpha}:\alpha\in\mathrm{Ord}\}$ by induction on $\alpha$ as follows: suppose we have already built $S$ up to some ordinal $\gamma$ as $\{s_{\delta}:\delta<\gamma\}$ such that:
$s_{\delta}<s_{\delta^{\prime}}$ and $f(s_{\delta})<f(s_{\delta^{\prime}})$ whenever $\delta<\delta^{\prime}<\gamma$,
and
$s_{\delta}\in Y_{\delta}$ for each $\delta<\gamma.$
@@
Let $\theta=\sup\{f(s_{\delta}):\delta<\gamma\}.$ Note that $\theta$ is well-defined because $\mathcal{M}$ satisfies the replacement scheme. Since we know that case 2 holds there must exist some $X$ in $\mathcal{F}_{n+1,\alpha}^{\ast}$ such that:
$f_{n+1,\alpha}^{-1}\{\beta:\beta<\gamma\}\cap X\neq\varnothing$,
therefore we can let $s_{\gamma\text{ }}$denote the first member of $X\cap Y_{\gamma\text{ }}$that is above $\{s_{\delta}:\delta<\gamma\}$. By conditions $(ii)$ and $(iv),$ $X\cap Y_{\gamma\text{ }}$is unbounded in the ordinals hence making it possible to define $s_{\gamma}$. Having defined the class $S$, let $X_{n+1,\alpha}=S$.
We are now done with the construction of $\left\{X_{n+1,\alpha}:\alpha\in\mathrm{Ord}^{\mathcal{M}}\right\}$. We pause to note that even though $\mathcal{U}$ itself in not definable in $\mathcal{M}$, for each fixed $n\in\omega$, the family $\mathcal{U}_{n}=\{X_{n,\alpha}:\alpha\in\mathrm{Ord}^{\mathcal{M}}\}$ is definable in $\mathcal{M}$, which is exactly why $\mathcal{U}$ satisfies condition (3) promised at the beginning of the proof. As in the Łoś proof of the fundamental theorem of ultraproducts (and thanks to the availability of an $\mathcal{M}$-definable global well-ordering) it is routine to check that the definable ultrapower:
$\mathcal{N}=\mathcal{M}^{\mathrm{Ord}^{\mathcal{M}}}/\ \mathcal{U}$,
consisting of $\mathcal{U}$-equivalence classes of $\mathcal{M}$-definable functions from $\mathrm{Ord}^{\mathcal{M}}$ to $M$ forms an elementary extension of $\mathcal{M}$. Also the construction makes it clear that $\mathcal{N}$ is taller than $\mathcal{M}$ since the equivalence class $[\mathrm{id}]_{U}$ exceeds all ordinals of $\mathcal{M}$ (where $\mathrm{id}$ is the identity map on $\mathrm{Ord}^{\mathcal{M}})$, thanks to the fact that $\mathcal{U}_{-1}\subseteq\mathcal{U}$. To verify conservativity of $\mathcal{N}$ over $\mathcal{M}$, suppose we are given a formula $\varphi(x,y)$ and some parameter $s\in N$, and we wish to show that the set $\{m\in M:\mathcal{N}\models\varphi(m,s)\}$ is $\mathcal{M}$-definable. Choose $f$ so that $s=$ the $\mathcal{U}$-equivalence class of $f=[f]_{\mathcal{U}}$ and choose a natural number $j$ large enough so that the formula $\theta(x,z)=\varphi(x,f(z))$ is a $\Sigma_{j}$-formula. Note that for each $m\in M$ the characteristic function of $S_{m}=\{z:\theta(m,z)\}$ is definable by some $\Sigma_{j}$-function. Hence for every $m\in M$ we have:
$S_{m}\in\mathcal{U}$ iff $S_{m}\in\mathcal{U}_{j},$
Therefore, if $\psi(\alpha,\beta)$ denotes the formula defining $\mathcal{U}_{j}$ in $\mathcal{M}$, then for all $m\in M,$ the following holds:
$\mathcal{N}\models\varphi(m,s)$ iff $\mathcal{M}\models\exists\alpha\forall\beta(\psi_{j}(\alpha,\beta)\rightarrow\gamma(\alpha,\beta)).$
@@
3.2. Corollary. Work in $\mathrm{ZF}+\exists p\left(\mathrm{V}=\mathrm{HOD}(p)\right)$ in the metatheory, and let $\mathcal{M}$ be an arbitrary model of $\mathrm{ZF}+\exists p\left(\mathrm{V}=\mathrm{HOD}(p)\right).$
(a) For every regular cardinal $\kappa$ there is a conservative elementary extension $\mathcal{N}_{\kappa}$ of $\mathcal{M}$ that is taller than $\mathcal{M}$ and $\mathrm{cf}(\mathrm{Ord}^{\mathcal{N}})=\kappa$ [^12] As indicated by the proof of this part, $\mathrm{ZFC}$ suffices in the metatheory.
(b) There is a proper class model $\mathcal{N}_{\mathrm{Ord}}$ whose elementary diagram is a definable class, $\mathcal{N}_{\mathrm{Ord}}$ is a conservative elementary extension of $\mathcal{M}$ that is taller than $\mathcal{M}$, and $\mathrm{Ord}^{\mathcal{N}_{\mathrm{Ord}}}$ contains a cofinal class of order-type $\mathrm{Ord}$.
@@
With part (a) at hand, it is evident that in the presence of a global definable well-ordering (available thanks to the veracity of $\exists p\left(\mathrm{V}=\mathrm{HOD}(p)\right)$ in the metatheory), we can construct the desired model $\mathcal{N}_{\mathrm{Ord}}$ as the union of an elementary chain of models $\left\langle\mathcal{M}_{\alpha}:\alpha\in\mathrm{Ord}\right\rangle$ that satisfies (1) and it also satisfies the result of replacing $\kappa$ with Ord in (2) and (3). $\square$
@@
3.4. Theorem. Every model of $\mathrm{ZF}+\exists p\left(\mathrm{V}=\mathrm{HOD}(p)\right)$ has a cofinal conservative elementary extension that possesses a minimal elementary end extension.
Proof. Let $\mathcal{M}$ be a model of $\mathrm{ZF}+\exists p\left(\mathrm{V}=\mathrm{HOD}(p)\right)$, $\mathcal{N}$ be the model constructed in the second proof of Theorem 3.1, and let $\mathcal{M}^{\ast}$ be the convex hull of $\mathcal{M}$ in $\mathcal{N}$. Note that by Gaifman Splitting Theorem we have:
@@
The above together with the fact that $\mathcal{N}$ is conservative extension of $\mathcal{M}$ makes it clear that $\mathcal{M}^{\ast}$ is a conservative extension of $\mathcal{M}$ [^14] By Theorem 4.1, $\mathcal{M}^{\ast}$ is a proper elementary extension of $\mathcal{M}$. We claim that $\mathcal{N}$ is a minimal elementary end extension of $\mathcal{M}^{\ast}$. First note that $\mathcal{N}$ is generated by members of $\mathcal{M}$ and the equivalence class of the identity function $[\mathrm{id}]_{\mathcal{U}}$ via the definable terms of $\mathcal{M}$, i.e., the following holds:
@@
For any $m\in M$, there exists some $\mathcal{M}$-definable function $f$ such that $\mathcal{N}\models f([\mathrm{id}]_{\mathcal{U}})=m.$
@@
To establish our claim it is sufficient to show that if $s\in N\backslash M^{\ast}$ then $[\mathrm{id}]_{\mathcal{U}}$ is definable in $\mathcal{N}$ using parameters from $M\cup\{s\}.$ Choose $h$ such that $s=[h]_{\mathcal{U}}$ and recall that by our construction of $\mathcal{U}$ we know that there exists $X\subseteq Y$ with $X\in\mathcal{U}$ such that either $h\upharpoonright X$ is one-to-one, or the image of $X$ under $h$ is a set (and thus of bounded rank). However, since $s$ was chosen to be in $N\backslash M^{\ast}$ we can rule out the latter possibility. Note that since $X\in\mathcal{U}$ the extension of $X$ in $\mathcal{N}$ will contain $[\mathrm{id}]_{\mathcal{U}}$, i.e., if $X$ is defined in $\mathcal{M}$ by $\varphi(x)$ then $\mathcal{N}\models\varphi([\mathrm{id}]_{\mathcal{U}})$. Moreover, by elementarity, since $f$ is one-to-one on $X$ in $\mathcal{M}$, it must remain so in $\mathcal{N}$. Hence we can define $[\mathrm{id}]_{\mathcal{U}}$ to be the unique element of $h^{-1}(b)\cap X.$ This completes the proof. $\square$
@@
3.5. Remark. As Theorem 4.1 of the next section shows, in Theorem 3.1 the model $\mathcal{N}$ cannot be required to fix $\omega^{\mathcal{M}}$ and in particular $\mathcal{N}$ cannot be arranged to end extend $\mathcal{M}$. By putting this observation together with Remark 2.6(k) we can conclude that in the proof of Corollary 3.2, $\omega^{\mathcal{N}_{\kappa}}$ is $\kappa$-like (i.e., its cardinality is $\kappa$ but each proper initial segment of it is of cardinality less than $\kappa$), and $\omega^{\mathcal{N}_{\mathrm{Ord}}}$ is Ord-like (i.e., it is a proper class every proper initial segment of which is isomorphic to an ordinal).
@@
We next address the question of the extent to which the hypothesis in Theorem 3.1 that $\mathcal{M}$ is a model of $\exists p\left(\mathrm{V}=\mathrm{HOD}(p)\right)$ can be wakened. In what follows $\mathrm{AC}$ stands for the axiom of choice, and $\mathrm{DC}$ stands for the axiom of dependent choice.
@@
3.6. Proposition. Suppose $\mathcal{M}$ and $\mathcal{N}$ are models of $\mathrm{ZF}$, $\mathcal{M}\prec_{\mathrm{cons}}\mathcal{N}$, and $\mathcal{N}$ is taller than $\mathcal{M}$. If every set can be linearly ordered in $\mathcal{N}$, then there is an $\mathcal{M}$-definable global linear ordering.
@@
Proof. Suppose $\mathcal{M}$ and $\mathcal{N}$ are as in the assumptions of the proposition. Fix $\alpha\in\mathrm{Ord}^{\mathcal{N}}$ such that $\alpha$ is above all the ordinals of $\mathcal{M}$ and let $R$ denote a linear ordering of $\mathrm{V}_{\alpha}^{\mathcal{N}}$ in $\mathcal{N}$. By conservativity, there is a formula $r(x,y)$ (possibly with parameters from $M$) such that $r(x,y)$ that defines $R\cap M^{2}$ in $\mathcal{M}$ It is clear that $r(x,y)$ defines a global linear ordering in $\mathcal{M}$. $\square$
@@
3.7. Corollary. There are models of $\mathrm{ZFC}$ that have no taller conservative elementary extensions to a model of $\mathrm{ZFC}$.
@@
Proof. This follows from Proposition 3.6, and the well-known fact that there are models of ZFC in which the universe is not definably linearly ordered. [^15] Easton proved (in his unpublished dissertation [Ea]) that assuming Con($\mathrm{ZF}$) there is a model $\mathcal{M}$ of $\mathrm{ZFC}$ that carries no $\mathcal{M}$-definable global choice function for the class of pairs in $\mathcal{M}$; and in particular the universe cannot be definably linearly ordered in $\mathcal{M}$. Easton’s theorem was exposited by Felgner [F, p.231]; for a more recent and streamlined account, see Hamkins’ MathOverflow answer [H]. $\square$
@@
3.9. Remark. If $\mathcal{M}\models\mathrm{ZF}$ and $\mathcal{M}\prec_{\Delta_{0},\ \mathrm{cons}}\mathcal{N}$, then each $k\in\omega^{\mathcal{M}}$ is fixed. To see this, suppose to the contrary that $r\in\mathrm{Ext}_{\mathcal{N}}(k)\backslash\mathrm{Ext}_{\mathcal{M}}(k)$. Consider the set $S=\left\{x\in\mathrm{Ext}_{\mathcal{M}}(k):\mathcal{N}\models x<r\right\}.$ It is easy to see that $S$ has no last element. On the other hand, by the assumption $\mathcal{M}\prec_{\mathrm{cons}}\mathcal{N}$, the set $S=$$\{x\in\mathrm{Ext}_{\mathcal{M}}(k):\mathcal{N}\models x<r\}$ is a definable subset of the predecessors of $r$ and thus by the veracity of the separation scheme in $\mathcal{M}$, $S$ is coded in $\mathcal{M}$ by some element, which makes it clear that $S$ must have a last element, contradiction. This shows that in Theorem 4.1, $\mathcal{N}$ fixes each $k\in\omega^{M}.$ A natural question is whether a taller conservative elementary extension of $\mathcal{M}$ can be arranged to fix $\omega^{\mathcal{M}}$. As we shall see in the next section, this question has a negative answer for models of $\mathrm{ZFC}$ (see Corollary 4.2).
4. Remark. Theorem 4.1 should also be contrasted with Gaifman’s theorem mentioned in Remark 2.5 which bars taller elementary self-extensions of models of (a fragment of) $\mathrm{ZF}$. Thus in Gaifman’s theorem, (a) the condition “$\mathcal{N}$ is an elementary self extension of $\mathcal{M}$” cannot be weakened to “$\mathcal{N}$ is an elementary conservative extension of $\mathcal{M}$”. Also note that in the first proof of Theorem 3.1, each $\mathcal{N}_{k}$ is a self-extension of $\mathcal{M}$ whose associated embedding $j$ is a $\Sigma_{k+2}$-elementary embedding whose image is not cofinal in $\mathcal{M}$.

### v6

@@ 3. $\mathbb{C}_{3}(n):$
If case 1 holds, then we let $P:=Y_{\gamma_{0}}$, where $\gamma_{0}$ is the first ordinal in $\mathrm{Ord}^{\mathcal{M}}$ witnessing the veracity of case 1. Clearly the range of $f\upharpoonright$ $P$ is coded as an element of $\mathcal{M}$.
If case 2 holds, then let $\mathfrak{X}^{\ast}:=$ the set of all $\mathcal{M}$-finite intersections of $\mathfrak{X}$. Note that $\mathfrak{X}^{\ast}$ is $\mathcal{M}$-listable, so we can let $\left\langle D_{\eta}:\eta\in\mathrm{Ord}^{\mathcal{M}}\right\rangle$ be an $\mathcal{M}$-list of $\mathfrak{X}^{\ast}$. It is also clear that the $\mathcal{M}$-thickness of $\mathfrak{X}$ is inherited by $\mathfrak{X}^{\ast}$. We will construct an $\mathcal{M}$-definable set $P=\{p_{\eta}:\eta\in\mathrm{Ord}^{\mathcal{M}}\}$ by recursion (within $\mathcal{M}$) such that the restriction of $f$ to $P$ is one-to-one, and such that $\mathfrak{X}\cup\{P\}$ is $\mathcal{M}$-thick. Suppose we have already built $P$ up to some $\gamma\in\mathrm{Ord}^{\mathcal{M}}$ as $\{p_{\eta}:\eta<\gamma\}$ such that the following two conditions hold in $\mathcal{M}$:
$(\ast)$ $p_{\eta}<p_{\eta^{\prime}}$ and $f(p_{\eta})<f(p_{\eta^{\prime}})$ whenever $\eta<\eta^{\prime}<\gamma$.
$(\ast\ast)$ $p_{\eta}\in$ $D_{\eta}$ for all $\eta<\gamma.$
@@
Within $\mathcal{M}$, let $\theta=\sup\{f(p_{\eta}):\eta<\gamma\}.$ Note that $\theta$ is well-defined because $\mathcal{M}$ satisfies the replacement scheme. Since we know that case 2 holds, there must exist some $D_{\delta}\in\mathfrak{X}^{\ast}$ such that:
$D_{\delta}\cap Y_{\theta+1}=\varnothing$, where $Y_{\theta+1}:=\{\alpha\in\mathrm{Ord}^{\mathcal{M}}:\mathcal{M}\models f(\alpha)<\theta+1\}.$
Therefore within $\mathcal{M}$ can choose $p_{\gamma}$ to be the first member of $D_{\delta}\cap D_{\gamma\text{ }}$that is above $\{s_{\eta}:\eta<\gamma\}$. Since $\mathfrak{X}$ is $\mathcal{M}$-thick and $\mathcal{F}\subseteq\mathfrak{X}$, $D_{\delta}\cap D_{\gamma\text{ }}$is unbounded in $\mathrm{Ord}^{\mathcal{M}},$ hence $s_{\gamma}$ is well-defined. Thus we have constructed an $\mathcal{M}$-definable set $P=\{p_{\eta}:\eta\in\mathrm{Ord}^{\mathcal{M}}\}$ such that $p_{\eta}\in$ $D_{\eta}$ for all $\eta\in\mathrm{Ord}^{\mathcal{M}}$ (so $\mathfrak{X}\cup\{P\}$ is $\mathcal{M}$-thick), and $f$ is strictly monotone increasing on $P$. This concludes the proof of part (b) of Lemma 3.1.1. $\square$
Using Lemma 3.1.1, it is now straightforward to use transfinite recursion within $\mathcal{M}$ of length $\mathrm{Ord}^{\mathcal{M}}$ to construct a family $\mathcal{U}_{n+1}\subseteq\mathbb{B}$ that extends $\mathcal{U}_{n}$ and which satisfies clauses $\mathbb{C}_{1}(n+1)$ through $\mathbb{C}_{3}(n+1)$. $\mathcal{U}_{n+1}$ will be of the form:
$\mathcal{U}_{n}\ \cup\bigcup\limits_{\alpha\in\mathrm{Ord}^{\mathcal{M}}}\{A_{\alpha},P_{\alpha}\}$,
where $A_{\alpha}$ and $P_{\alpha}$ are defined within $\mathcal{M}$ by transfinite recursion on $\alpha$. Suppose for some $\alpha\in\mathrm{Ord}^{\mathcal{M}}$ we have built subsets $A_{\beta}$ and $P_{\beta}$ of $\mathbb{B}$ for each $\beta<\alpha,$ and $\mathcal{U}_{n}\cup\bigcup\limits_{\beta<\alpha}\{A_{\beta},P_{\beta}\}$ is $\mathcal{M}$-thick. Now we can use part (a) of Lemma 3.1.1 to construct $A_{\alpha}\in\mathbb{B}$ such that $\mathcal{U}_{n}\cup\bigcup\limits_{\beta<\alpha}\{A_{\beta},P_{\beta}\}\cup\{A_{\alpha}\}$ is $\mathcal{M}$-thick and $A_{\alpha}=S_{n,\alpha}$ or $A_{\alpha}=\mathrm{Ord}^{\mathcal{M}}\backslash S_{n,\alpha}.$ In the next step, using part (b) of Lemma 3.1.1, we can construct $P_{\alpha}\in\mathbb{B}$ which ‘takes care of’ $f_{n,\alpha}$ (i.e., the restriction of $f_{n,\alpha}$ to $P_{\alpha}$ is either one-to-one, or its range is a set) and which has the property that $\mathcal{U}_{n}\cup\bigcup\limits_{\beta<\alpha}\{A_{\beta},P_{\beta}\}\cup\{A_{\alpha},P_{\alpha}\}$ is $\mathcal{M}$-thick. This concludes the recursive construction of $A_{\alpha}$ and $P_{\alpha}$ for $\alpha\in\mathrm{Ord}^{\mathcal{M}}$, thus we can define:
$\mathcal{U}_{n+1}:=\mathcal{U}_{n}\ \cup\bigcup\limits_{\alpha\in\mathrm{Ord}^{\mathcal{M}}}\{A_{\alpha},P_{\alpha}\}$.
The construction makes it clear that $\mathcal{U}_{n+1}$ satisfies $\mathbb{C}_{1}(n+1)$ through $\mathbb{C}_{3}(n+1)$. This, in turn, makes evident that for
$\mathcal{U}:=\bigcup\limits_{n\in\omega}\mathcal{U}_{n}$,
$\mathcal{U}$ has properties (1) through (4) promised at the beginning of Stage 1. Note that $\mathcal{U}$ satisfies condition (4) since $\mathcal{U}_{n+1}$ and $\mathbb{B}_{n}$ are both $\mathcal{M}$-listable, and by clause $\mathbb{C}_{3}(n+1)$, $\mathcal{U}\cap\mathbb{B}_{n}\subseteq\mathcal{U}_{n+1}.$ This concludes Stage 1.
STAGE 2: Let $\mathcal{N}:=\mathcal{M}^{\mathrm{Ord}^{\mathcal{M}}}/\ \mathcal{U}$ be the definable ultrapower of $\mathcal{M}$ modulo $\mathcal{U}$ whose universe consists of $\mathcal{U}$-equivalence classes $f]_{\mathcal{U}}$ of $\mathcal{M}$-definable functions $f:\mathrm{Ord}^{\mathcal{M}}\rightarrow M$. Thanks to the availability of an $\mathcal{M}$-definable global well-ordering, the following Łoś-style result can be readily verified.
Theorem 3.1.2. For all $k$-ary formulae $\varphi(x_{0},\cdot\cdot\cdot,x_{k-1})$ of $\mathcal{L}_{\mathrm{set}}$ and all $k$-tuples of $\mathcal{M}$-definable functions $\left\langle f_{i}:i<k\right\rangle$ from $\mathrm{Ord}^{\mathcal{M}}$ to $M$, we have:
$\mathcal{N}\models\varphi([f_{0}]_{\mathcal{U}},\cdot\cdot\cdot,[f_{k-1}]_{\mathcal{U}})$ iff $\left\{\alpha\in\mathrm{Ord}^{\mathcal{M}}:\mathcal{M}\models\varphi\left(f_{0}(\alpha),\cdot\cdot\cdot,f_{k-1}(\alpha)\right)\right\}\in\mathcal{U}$.
For each $m\in M$ let $\widetilde{m}$: $\mathrm{Ord}^{\mathcal{M}}\rightarrow\{m\}$. Theorem 3.1.2 assures us that the map $j:\mathcal{M}\rightarrow\mathcal{N}$ given by $j(m)=[\widetilde{m}]_{\mathcal{U}}$ is an elementary embedding, thus by identifying $\mathcal{M}$ with the image of $j$, we can construe $\mathcal{N}$ as an elementary extension of $\mathcal{M}$. Let $\mathrm{id}:\mathrm{Ord}^{\mathcal{M}}\rightarrow\mathrm{Ord}^{\mathcal{M}}$ be the identity map. By property (2) of $\mathcal{U}$ specified in Stage 1, every element of $\mathcal{U}$ is unbounded in $\mathrm{Ord}^{\mathcal{M}}.$ So by Theorem 3.1.2, $\mathcal{N}$ is taller than $\mathcal{M}$ since $\mathrm{id}]_{\mathcal{U}}$ exceeds all ordinals of $\mathcal{M}$.
It remains to verify that $\mathcal{N}$ is a conservative extension of $\mathcal{M}$. Given a formula $\psi(x,y_{0},\cdot\cdot\cdot,y_{k-1})$ and parameters $s_{0},\cdot\cdot\cdot,s_{k-1}$ in $N$, let:
$Y=\left\{m\in M:\mathcal{N}\models\varphi([\widetilde{m}]_{\mathcal{U}},s_{0},\cdot\cdot\cdot,s_{k-1})\right\}.$
We wish to show that $Y$ is $\mathcal{M}$-definable. Choose $f_{0},\cdot\cdot\cdot,f_{k-1}$ so that $s_{i}=[f_{i}]_{\mathcal{U}}$ for all $i<p$ and consider the formula $\theta(x,\alpha)$ defined below:
$\theta(x,\alpha):=\alpha\in\mathrm{Ord}\wedge\varphi(x,f_{0}(\alpha),\cdot\cdot\cdot,f_{k-1}(\alpha))$.
Pick $n\in\omega$ large enough so that $\theta(x,\alpha)$ is a $\Sigma_{n}$-formula. For each $m\in M,$ let:
$S_{m}:=\{\alpha\in\mathrm{Ord}^{\mathcal{M}}:\mathcal{M}\models\theta(m,\alpha)\}$.
Note that the choice of $n$, $S_{m}\in\mathbb{B}_{n}$ for each $m\in M$, so $S_{m}\in\mathcal{U}$ iff $S_{m}\in\mathcal{U}\cap\mathbb{B}_{n}$ for each $m\in M$. On the other hand, by Theorem 3.1.2 we have:
$m\in Y$ iff $S_{m}\in\mathcal{U}$ for each $m\in M$.
So we can conclude that $m\in Y$ iff $S_{m}\in\mathcal{U}\cap\mathbb{B}_{n}$ for each $m\in M.$ Since $\mathcal{U}\cap\mathbb{B}_{n}$ is $\mathcal{M}$-listable by property (4) of $\mathcal{U}$ specified in Stage 1, we can write $\mathcal{U}\cap\mathbb{B}_{n}$ as an $\mathcal{M}$-list $\left\langle X_{\eta}:\eta\in\mathrm{Ord}^{\mathcal{M}}\right\rangle$ given by some formula $\psi(x,y).$ Thus the following holds for all $m\in M$, which shows that $Y$ is $\mathcal{M}$-definable.
$m\in Y$ iff $\exists\eta\left(S_{m}=X_{\eta}\right)$ iff $\mathcal{M}\models\exists\eta\ \forall\alpha\left(\theta(m,\alpha)\leftrightarrow\psi(\eta,m)\right).$
@@
3.2. Corollary. Work in $\mathsf{ZF}+\exists p\left(\mathrm{V}=\mathrm{HOD}(p)\right)$ in the metatheory, and let $\mathcal{M}$ be an arbitrary model of $\mathsf{ZF}+\exists p\left(\mathrm{V}=\mathrm{HOD}(p)\right).$
  1. (a)
    For every regular cardinal $\kappa$ there is a conservative elementary extension $\mathcal{N}_{\kappa}$ of $\mathcal{M}$ that is taller than $\mathcal{M}$ and $\mathrm{cf}(\mathrm{Ord}^{\mathcal{N}})=\kappa$. [^22] As indicated by the proof of this part, $\mathsf{ZFC}$ suffices as the metatheory for this part.
  2. (b)
    There is a proper class model $\mathcal{N}_{\mathrm{Ord}}$ whose elementary diagram is a definable class such that $(1)$ $\mathcal{M\prec}_{\mathrm{cons,taller}}\mathcal{N}_{\mathrm{Ord}}$, and $(2)$ $\mathrm{Ord}^{\mathcal{N}_{\mathrm{Ord}}}$ contains a cofinal class of order-type $\mathrm{Ord}$.
@@
With part (a) at hand, it is evident that in the presence of a global definable well-ordering (available thanks to the veracity of $\exists p\left(\mathrm{V}=\mathrm{HOD}(p)\right)$ in the metatheory), we can construct the desired model $\mathcal{N}_{\mathrm{Ord}}$ as the union of an elementary chain of models $\left\langle\mathcal{M}_{\alpha}:\alpha\in\mathrm{Ord}\right\rangle$ that satisfies (1) and it also satisfies the result of replacing $\kappa$ with $\mathrm{Ord}$ in (2) and (3). $\square$
@@
3.4. Theorem. Every model of $\mathsf{ZF}+\exists p\left(\mathrm{V}=\mathrm{HOD}(p)\right)$ has a cofinal conservative elementary extension that possesses a minimal elementary end extension.
Proof. Let $\mathcal{M}$ be a model of $\mathsf{ZF}+\exists p\left(\mathrm{V}=\mathrm{HOD}(p)\right)$, $\mathcal{N}$ be the model constructed in the second proof of Theorem 3.1, and let $\mathcal{M}^{\ast}$ be the convex hull of $\mathcal{M}$ in $\mathcal{N}$. Note that by Gaifman Splitting Theorem we have:
@@
The above together with the fact that $\mathcal{N}$ is conservative extension of $\mathcal{M}$ makes it clear that $\mathcal{M}^{\ast}$ is a conservative extension of $\mathcal{M}$. [^24] By Theorem 4.1, $\mathcal{M}^{\ast}$ is a proper elementary extension of $\mathcal{M}$. We claim that $\mathcal{N}$ is a minimal elementary end extension of $\mathcal{M}^{\ast}$. First note that $\mathcal{N}$ is generated by members of $\mathcal{M}$ and the $\mathcal{U}$-equivalence class of the identity function $\mathrm{id}]_{\mathcal{U}}$ via the definable terms of $\mathcal{M}$, i.e., the following holds:
@@
For any $a\in N$, there exists some $\mathcal{M}$-definable function $f$ such that $\mathcal{N}\models f([\mathrm{id}]_{\mathcal{U}})=a.$
@@
To verify that $\mathcal{N}$ is a minimal elementary end extension of $\mathcal{M}^{\ast}$ it is sufficient to show that if $s\in N\backslash M^{\ast}$ then $\mathrm{id}]_{\mathcal{U}}$ is definable in $\mathcal{N}$ using parameters from $M\cup\{s\}.$ Choose $h$ such that $s=[h]_{\mathcal{U}}$ and recall that by our construction of $\mathcal{U}$ we know that there exists $X\subseteq\mathrm{Ord}^{\mathcal{M}}$ with $X\in\mathcal{U}$ such that either $h\upharpoonright X$ is one-to-one, or the image of $X$ under $h$ is a set (and thus of bounded $\mathcal{M}$-rank). However, since $s$ was chosen to be in $N\backslash M^{\ast}$ we can rule out the latter possibility. Note that since $X\in\mathcal{U}$ the extension of $X$ in $\mathcal{N}$ will contain $\mathrm{id}]_{\mathcal{U}}$, i.e., if $X$ is defined in $\mathcal{M}$ by $\varphi(x)$, then $\mathcal{N}\models\varphi([\mathrm{id}]_{\mathcal{U}})$. Moreover, by elementarity, since $h$ is one-to-one on $X$ in $\mathcal{M}$, it must remain so in $\mathcal{N}$. Hence, in light of the fact that $\mathcal{N}\models h([\mathrm{id}]_{\mathcal{U}})=s$, we can define $\mathrm{id}]_{\mathcal{U}}$ within $\mathcal{N}$ to be the unique element of $h^{-1}(s)\cap X.$ This completes the proof. $\square$
@@
3.5. Remark. As Theorem 4.1 of the next section shows, in Theorem 3.1 the model $\mathcal{N}$ cannot be required to fix $\omega^{\mathcal{M}}$ and in particular $\mathcal{N}$ cannot be arranged to end extend $\mathcal{M}$. By putting this observation together with Remark 2.6(j) we can conclude that in the proof of Corollary 3.2, $\omega^{\mathcal{N}_{\kappa}}$ is $\kappa$-like (i.e., its cardinality is $\kappa$ but each proper initial segment of it is of cardinality less than $\kappa$), and $\omega^{\mathcal{N}_{\mathrm{Ord}}}$ is $\mathrm{Ord}$-like (i.e., it is a proper class every proper initial segment of which forms a set).
@@
We next address the question of the extent to which the hypothesis in Theorem 3.1 that $\mathcal{M}$ is a model of $\exists p\left(\mathrm{V}=\mathrm{HOD}(p)\right)$ can be weakened. In what follows $\mathrm{AC}$ stands for the axiom of choice, and $\mathsf{DC}$ stands for the axiom of dependent choice.
@@
3.6. Proposition. Suppose $\mathcal{M}$ and $\mathcal{N}$ are models of $\mathsf{ZF}$, $\mathcal{M}\prec_{\mathrm{cons}}\mathcal{N}$, and $\mathcal{N}$ is taller than $\mathcal{M}$. If every set can be linearly ordered in $\mathcal{N}$, then there is an $\mathcal{M}$-definable global linear ordering.
@@
Proof. Suppose $\mathcal{M}$ and $\mathcal{N}$ are as in the assumptions of the proposition. Fix $\alpha\in\mathrm{Ord}^{\mathcal{N}}$ such that $\alpha$ is above all the ordinals of $\mathcal{M}$ and let $R$ denote a linear ordering of $\mathrm{V}_{\alpha}^{\mathcal{N}}$ in $\mathcal{N}$. By conservativity, there is a formula $r(x,y)$ (possibly with parameters from $M$) such that $r(x,y)$ that defines $R\cap M^{2}$ in $\mathcal{M}$ . It is clear that $r(x,y)$ defines a global linear ordering in $\mathcal{M}$. $\square$
@@
3.7. Corollary. There are models of $\mathsf{ZFC}$ that have no taller conservative elementary extensions to a model of $\mathsf{ZFC}$.
@@
Proof. This follows from Proposition 3.6, and the well-known fact that there are models of $\mathsf{ZFC}$ in which the universe is not definably linearly ordered. [^25] Easton proved (in his unpublished dissertation [Ea]) that assuming $\mathrm{Con}(\mathsf{ZF})$ there is a model $\mathcal{M}$ of $\mathsf{ZFC}$ that carries no $\mathcal{M}$-definable global choice function for the class of pairs in $\mathcal{M}$; and in particular the universe cannot be definably linearly ordered in $\mathcal{M}$. Easton’s theorem was exposited by Felgner [F, p.231]; for a more recent and streamlined account, see Hamkins’ MathOverflow answer [H]. $\square$
@@
[…]
3.10. Remark. Theorem 4.1 should also be contrasted with Gaifman’s theorem mentioned in Remark 2.5 which bars taller elementary self-extensions of models of (a fragment of) $\mathsf{ZF}$. Thus in Gaifman’s theorem, (a) the condition $\mathcal{N}$ is an elementary self extension of $\mathcal{M}$” cannot be weakened to $\mathcal{N}$ is an elementary conservative extension of $\mathcal{M}$”. Also note that in the first proof of Theorem 3.1, each $\mathcal{N}_{k}$ is a self-extension of $\mathcal{M}$ whose associated embedding $j$ is a $\Sigma_{k+2}$-elementary embedding whose image is not cofinal in $\mathcal{M}$.